OpenStudy (anonymous):

Suppose that g is defined as g(x) = -6(x-2)^2-2 what is the range?

OpenStudy (asnaseer):

this site might help you understand range and its related concepts of domain and co-domain: http://www.mathsisfun.com/sets/domain-range-codomain.html

OpenStudy (anonymous):

yes i understand what domain and range are i just dont understand how you connect it to a number line??

OpenStudy (asnaseer):

look at the function you were given:\[g(x)=-6(x-2)^2-2\]notice the term \((x-2)^2\) will ALWAYS be positive, so its value will range from 0 to \(\infty\), agreed?

OpenStudy (anonymous):

yes so its not range is less than or equal -2

OpenStudy (asnaseer):

so maximum value of g(x) will be when \((x-2)^2=0\) which gives g(x)=-2. so g(x) ranges from -2 to -infinity

OpenStudy (asnaseer):

\[-\infty<g(x)\le-2\]

OpenStudy (anonymous):

so on number line i would put a closed circle on -2 and then have a line go infinitely onto the negative direction

OpenStudy (asnaseer):

precisely - you got it!

OpenStudy (anonymous):

thnks

OpenStudy (asnaseer):

yw :)